Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name jpchen
Log entry time 20:40:00 on October 8,2010
Entry number 333957
keyword=Loop 1 is D2 and is already cold now.
At around 7 pm, ESR 15 K flow to Hall A increased from 6g/s to 14g/s.
Loop 1 cool-down started automatically.
Around 7:30 pm TO (Bo) called me that there is a low-pressure alarm on D2 gas panel monitor. At that time, the flow was >20g/s. Appearantly we already got loop 1 at 19K, but H2 pressure is still high (~45psi) while D2 pressure is low ~18-19 psi. We reduced flow to ~10 g/s. I came to the counting house checked all the readings and verified in the hall. It became clear that D2 gas is in loop 1 (not loop2) and H2 gas is in loop 2.
I re-adjusted the temperature settings to have loop 1 (D2) at 22K.
Set fan speek to 36 HZ. It is ready.
We are now start to also cool down loop 2 (H2).